The velocity of sound in air at 4 atmosphere and that at 1 atmosphere pressure would be
(a) 1 : 1
(b) 4 : 1
(c) 1 : 4
(d) 3 : 1

Grade:Upto college level

2 Answers

vaibhav patil
3 years ago
the velocity of sound in  air is depends on tempereture and 
there is no effect of pressure

\because \frac{P}{\rho }= \frac{RT}{M}

\therefore V= \sqrt{\frac{\gamma RT}{M}}

 

- wherein

V \alpha \sqrt{T}

T is in Kelvin

So velocity increses with increase in temperature.

 

 

 Since V=\sqrt{\frac{\gamma P}{\rho}}=\sqrt{\frac{\gamma RT}{M}}

V is independent of P

\therefore Ratio of velocity = 1:1

answer

A

Vikas TU
3 years ago
The ratio is 1. Speed of sound depends only on temperature (and the properties of the gas). It doesn't depend on pressure.